OpenStudy (anonymous):

One number is eight more than twice another. If their difference is 25, what is the larger number? 17 42 35

OpenStudy (anonymous):

Two numbers a and b a=2b+8 a=b+25 b+25= 2b+8 b+17=2b b=17 a+8=34+8=42 a=42 42